# PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis

### v1

@@ B.1 Performance breakdown: monocular setting
B.2 Performance breakdown: multiview setting
We provide per-scene breakdown performance for the multiview video reconstruction setting in Table 10. We report results for D-3DGS [36], 4DGS [37], and STG [38], which serve as the most common and recent baselines for the DyNVS task. Note that since SOM [15] is specialized for monocular video setups, we omit its performance in the multiview evaluation.
Table 10: Per-scene breakdown results for all 17 scenes under the multiview setting.
@@ B.2 Performance breakdown: multiview setting
B.3 Additional Qualitative Results
Figure 10 shows additional qualitative results, where all methods fail to accurately capture multi-body interactions, producing blurred or under-reconstructed outputs. For the 4DGS [37], using the default grid learning rate frequently causes NaN values in the loss. Reducing the learning rate stabilizes training but results in poor dynamic capture, with outputs resembling static scenes.
@@ B.3 Additional Qualitative Results
(e) SOM [15]
Figure 10: Qualitative results of recent DyNVS methods on the basin, bouncing balls, box-smoke, pancake, and ice scenes with monocular training setup. These results show that all methods frequently exhibit needle-like artifacts and fail to reconstruct dynamic elements accurately.
Appendix C License
Our dataset is released under the Creative Commons Attribution-NonCommercial (CC BY-NC) license. The mesh used for the lucy scene is sourced from the Stanford 3D Repository, which permits usage for research purposes. Other mesh objects are licensed under CC-BY 4.0, as detailed in Table 11. For the texture maps in the textile categories, we utilized images from Pixabay, which are freely available with contributor consent and comparable to a CC-BY 4.0 license, also summarized in Table 11. Therefore, all data included in our dataset comply with usage rights and do not pose intellectual property issues.
Table 11: License of the sources used for generating data.
@@ Appendix C License
Appendix D Limitations & Broad Impact
Limitations
Although our PhysGaia dataset holds a crucial position in the DyNVS field by enabling accurate evaluation with ground-truth physical properties, it is inherently limited to synthetic scenarios. To bridge this gap towards more realistic applications, we believe domain adaptation approaches, such as those from Syn2Real [79], could be adopted. However, such methods must be carefully designed to maintain multi-view consistency across frames. We leave this important domain transfer for future work, confident that our dataset’s precise physical information will be invaluable for developing these advanced techniques.
Broad impact

### v2

@@ E.3 Performance breakdown: monocular setting
E.4 Performance breakdown: multiview setting
We provide per-scene breakdown performance for the multiview video reconstruction setting in Table G. We report results for D-3DGS [79], 4DGS [74], and STG [42], which serve as the most common and recent baselines for the DyNVS task. Note that since MoSca [38] and SOM [70] are specialized for monocular video setups, we omit its performance in the multiview evaluation.
Table G: Per-scene breakdown results for all 17 scenes under the multiview setting.
@@ E.4 Performance breakdown: multiview setting
E.5 Additional Qualitative Results
Figure D and Figure E present qualitative results from monocular settings, revealing that all methods struggle to accurately capture multi-body interactions. This leads to common issues like needle-like artifacts and under-reconstruction of dynamic regions. For 4DGS [74], its default grid learning rate often causes NaN loss values; while reducing it stabilizes training, the output resembles static scenes due to poor dynamic capture.
@@ E.5 Additional Qualitative Results
(e) MoSca [38]
(f) SOM [70]
Figure D: Qualitative results of recent DyNVS methods on the basin, bouncing balls, box-smoke, pancake, and ice scenes with monocular training setup. These results show that all methods frequently exhibit needle-like artifacts and fail to reconstruct dynamic elements accurately.
(a) GT image
(b) D-3DGS [79]
(c) 4DGS [74]
(d) STG [42]
(e) MoSca (2024)
(f) SOM [70]
Figure E: Qualitative results of recent DyNVS methods on the lucy, and hanok scenes with monocular setup. All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
F Limitations & Broad Impact
Limitations
While our PhysGaia is built upon a physics simulator to provide physically consistent supervision, it inherits the inherent approximations of the underlying physical models. In practice, physics simulators rely on simplified constitutive assumptions and numerical discretizations to make the problem tractable and stable. As a result, certain aspects of real-world dynamics, such as complex material responses or fine-scale interactions, may not be fully captured. We view this not as a limitation of our dataset alone, but as an inherent characteristic of simulation-based approaches, and an important direction for future work.
Broad impact
G License
Our benchmark is released under the Creative Commons Attribution-NonCommercial (CC BY-NC) license. The mesh used for the lucy scene is sourced from the Stanford 3D Repository, which permits usage for research purposes. Other mesh objects are licensed under CC-BY 4.0, as detailed in Table H. For the texture maps in the textile categories, we utilized images from Pixabay, which are freely available with contributor consent and comparable to a CC-BY 4.0 license, also summarized in Table H. Therefore, all data included in our benchmark comply with usage rights and do not pose intellectual property issues.
Table H: License of the sources used for generating data. The licenses for the mesh objects follow Creative Commons terms, except for the Lucy mesh sourced from the Stanford 3D Repository [37]. The images used in the PhysGaia scenes, along with their download sources, are licensed under the Content License granted by Pixabay.